# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 91             Set $Move=0$
107                   Declare $Current-Node$ as the black hole node and continue perpetual exploration avoiding this node.
@@
125             Declare the next node along the direction component of dir as the black hole node and continue perpetual exploration avoiding this node.
@@
5.2.1 Correctness and Complexity
In the next two lemmas (Lemma 5.9 and Lemma 5.11) we first ensure that when we say an agent encounters an anomaly, is actually an anomaly. And what an agents interpret from these anomalies are true. Next in Lemma 5.13 we ensure that the agent that changes to state Gather can always identify the state of the consumed agent at the time it gets destroyed at the black hole.
@@ 5.2.1 Correctness and Complexity
Let $a_{i}$ and $a_{j}$ be two agents exploring the segments $Seg(a_{i})$ and $Seg(a_{j})$, such that $Seg(a_{i})\cap Seg(a_{j})=v_{c}$, where $v_{c}$ is also the $home$ of $a_{j}$. If during execution of PerpExplore-Scat-Whitbrd, there exists a round $t>0$, in which $a_{i}$ in state Forward finds a visited type message, then there exists a round $0<t^{\prime}<t$, in which $a_{j}$ has been consumed by the black hole, while exploring the segment $Seg(a_{j})$ in state Backtrack.
Proof 5.10.
@@
Let us suppose, there does not exist any round $t^{\prime}<t$ in which $a_{j}$ has been consumed by the black hole in state Backtrack. This means that, either $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$ or $a_{j}$ has been consumed by the black hole at round $t^{\prime}$, while it is in state Forward.
@@
Case I: Let $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$. Note that since $a_{i}$ is at $v_{c}$ and checks for the visited type message, this implies $T_{time}=n$. This means, at $(t-n-1)$-th round $a_{i}$ changed its state to Forward from Initial. This also implies $a_{j}$ was also in state Initial at $(t-n-1)-$th round at $v_{c}$. Hence, at this round $a_{j}$ must have cleared any whiteboard data at $v_{c}$ (as $v_{c}$ is the $home$ of $a_{j}$) and also has moved clockwise after changing the state to Forward. Observe that, the node $v_{c}$ can only be explored by $a_{i}$ and $a_{j}$. So, these two agents has the possibility to write the visited type message at $v_{c}$ after round $t-n-1$ and before round $t$. Further, note that in the next $n+1$ rounds after $(t-n-1)-$th round (i.e., until round $t$), $a_{j}$ cannot return back to $v_{c}$. So, $a_{j}$ can not write anything on $v_{c}$ after round $t-n-1$ and before round $t$. Also, $a_{i}$ can only write the visited type message at $v_{c}$ at the $t-$th round (i.e., when $T_{time}=n$ according to Algorithm 4). But, in this case, $a_{i}$ at $t-$th round before writing a message, finds that there already exists a visited type message, and this a contradiction, to the fact that $a_{j}$ has not been consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$.
@@
Case II: Let $a_{j}$ has been consumed by the black hole at some round $t^{\prime}<t$, while in state Forward. This implies there exists a round $0<t^{\prime\prime}<t^{\prime}$ in which $a_{j}$ was in state Initial (this is the last time $a_{j}$ was in state Initial before it gets consumed by the black hole). Thus this means, $t^{\prime}\leq t^{\prime\prime}+n-2$ as a segment can be of length at most $n-2$. Then at round $t^{\prime\prime}$, $a_{i}$ was also in state Initial at its own $home$. Now, note that after $t^{\prime\prime}$, $a_{i}$ can visit $v_{c}$ within round $t^{\prime\prime}+1$ and $t^{\prime\prime}+n+1$ while in state Forward. Note that from round $t^{\prime\prime}+1$ and before round $t^{\prime\prime}+n+1$ no agent can write a visited type message at $v_{c}$ (by similar argument as in Case I). So during round $t^{\prime\prime}+n+1$, $a_{i}$ can not see any visited type message at $v_{c}$. So, $t>t^{\prime\prime}+n+1$. Also note that the next time after round $t^{\prime\prime}+n+1$, $a_{i}$ visits $v_{c}$ in state Forward, is earliest at round $t^{\prime\prime}+4n+1$. This implies $t\geq t^{\prime\prime}+4n+1$. Now, let us consider the agent $a_{k}$ where $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{k}$ and $Seg(a_{j})\cap Seg(a_{i})=home$ of $a_{i}$. Note that in round $t^{\prime\prime}+n+1$ both $a_{k}$ and $a_{i}$ moves into state Back-Wait (as none of them sees any visited type message). Next after waiting there upto round $t^{\prime\prime}+2n$, both of them changes their state to Backtrack at round $t^{\prime\prime}+2n+1$. Next they reach their corresponding $home$ and checks for a visited type message at round $t^{\prime\prime}+3n+1$. Note that $a_{i}$ finds the visited type message left by $a_{k}$ and changes to state Initial-Wait whereas, $a_{k}$ does not find any visited type message left by $a_{j}$ (as $a_{j}$ is consumed at the black hole at state Forward at round $t^{\prime}\leq t^{\prime\prime}+n-2<t^{\prime\prime}+3n+1$). So, $a_{k}$ changes to state Gather1 at round $t^{\prime\prime}+3n+1$ and moves clockwise with the message dir$=$ (Clockwise, NULL) until it finds $a_{i}$ at it’s $home$ in state Initial-Wait before round $t^{\prime\prime}+4n$ (as length of a segment can be of length at most $n-2$). Thus before $t^{\prime\prime}+4n+1$-th round $a_{i}$ changes its state to Gather1. Since no agent moves to state Forward from state Gather1, $a_{i}$ can not be at $v_{c}$ in state Forward during round $t$. So we again arrive at a contradiction. This implies $a_{j}$ can not be in state Forward while it is consumed at the black hole at round $t^{\prime}<t$.
@@
Let $a_{i}$ and $a_{j}$ be two agents such that $v_{c}=Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$. If there exists a round $t>0$ such that $a_{i}$ checks and finds no visited type message on $v_{c}$ while in state Backtrack, then there must be a round $t^{\prime}<t$ in which $a_{j}$ was consumed by the black hole while in state Forward.
Proof 5.12.
@@
Let us consider that there does not exists any round $t^{\prime}<t$ where $a_{j}$ was consumed by the black hole in state Forward. This implies that either $a_{j}$ is not consumed by the black hole for all rounds $t^{\prime}<t$ or it is consumed by the black hole at round $t^{\prime}<t$ while it was in state Backtrack.
@@
Case-I: Let us consider that $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime},\leavevmode\nobreak\ \forall t>t^{\prime}>0$. Note that, at round $t$, $a_{i}$ checks for a visited type message while in state Backtrack, and that means the current $T_{time}$ for any alive agent must be equal $3n$ (at $T_{time}=n$ every alive agent ends its state Forward, then until $T_{time}=2n$ every such agent is in state Back-Wait, and at $T_{time}=3n$ every alive agent checks for the message in state Backtrack). Note that, this means at time $t-3n-1$, both $a_{i}$ and $a_{j}$ was in state Initial at their respective $home$. Then at round $t-2n-1$, both of them must be in state Forward with $T_{time}=2n$, while $a_{j}$ is at $v_{c}$ and $a_{i}$ is at $home$ of $a_{k}$ ( $a_{k}$ is the third agent exploring $Seg(a_{k})$, such that $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$). So, at this point, both $a_{i}$ and $a_{j}$ must have written a visited type message at their respective current nodes (i.e., $v_{c}$ and $home$ of $a_{k}$). Further observe that, within $t-2n$ round to $t-n-1$ round $a_{j}$ remains at $v_{c}$ while in state Back-Wait, and according to our algorithm no agent in state Back-Wait can alter the already stored information at their current node, i.e., within these rounds $a_{j}$ cannot erase the visited type message at $v_{c}$. Next, within $t-n$ to $t-1$ round, the only agent which can visit the node $v_{c}$ is $a_{i}$, while it is in state Backtrack. Note that, $a_{i}$ cannot alter any visited type message during state Backtrack. Thus, $a_{i}$ at round $t$ checks and finds a visited type message at $v_{c}$. This is a contradiction to the assumption that $a_{i}$ does not find any visited type message at round $t$. Thus, $a_{j}$ must have been consumed by the black hole at time $t^{\prime}$ for some $t^{\prime}<t$.
@@
Case-II: Suppose $a_{j}$ has been consumed by the black hole at some round $t^{\prime}<t$, while in state Backtrack. Let $t^{\prime\prime}$ be the round when $a_{j}$ was in state Initial for the last time where $0<t^{\prime\prime}<t^{\prime}<t$. Note that $t^{\prime}>t^{\prime\prime}+2n+1$ (because an agent changes its state to Backtrack at round $t^{\prime\prime}+2n+1$). So at round $t^{\prime\prime}+n+1$ it was alive and at $v_{c}$ in state Forward. During this round it also writes a visited type message at $v_{c}$. Now as we discussed earlier in Case-I, from round $t^{\prime\prime}+n+1$ till round $t-1$ no agent can erase or alter the visited type message at $v_{c}$. Thus at round $t$, $a_{i}$ must find the visited type message at $v_{c}$ upon checking while in state Backtrack. This is a contradiction to the fact that at round $t$, $a_{i}$ checks and finds no visited type message at $v_{c}$ while it is in state Backtrack. Thus if $a_{j}$ is consumed by the black hole at round $t^{\prime}$, then it must be in state Forward.
@@
The agent which changes its state to Gather, correctly identifies the state in which an agent was in, just before it gets consumed by the black hole.
Proof 5.14.
@@
Let $a_{i},a_{j}$ and $a_{k}$ be three agents initially at three different nodes (i.e., three different $home$ for three agents) of $R$. Let $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$. Without loss of generality let $a_{j}$ be the agent that is consumed by the black hole at a round $t>0$. Now we have two cases according th the state of $a_{j}$ at the time of getting consumed.
@@
Case-I: $a_{j}$ gets consumed by the black hole while in state Forward at round $t$. Let $t^{\prime}<t$ be the round when $a_{j}$ was in state Initial for the last time. Note that at round $t^{\prime}$, all agents are in state Initial. We claim that $a_{i}$ is the agent to change its state to Gather. If possible let $a_{k}$ changes its state to Gather. Then first we show that it must change its state to Gather before round $t^{\prime}+3n+1$. Otherwise, since $a_{j}$, fails to reach $home$ of $a_{i}$ at round $t^{\prime}+n+1$, it does not write any visited type message there but $a_{i}$ reaches $home$ of $a_{k}$ at round $t^{\prime}+n+1$ and writes a visited type message at $home$ of $a_{k}$. Also, since, $a_{i}$ and $a_{j}$ are the only two agent which visits $home$ of $a_{i}$., this implies, no message can be written at $home$ of $a_{i}$ on and between rounds $t^{\prime}$ and $t^{\prime}+3n$. So, when $a_{i}$ returns $home$ and finds no visited type message round $t^{\prime}+3n+1$ while in state Backtrack, it changes its state to Gather but $a_{k}$ doesn’t do so as it sees the visited type message at its $home$, left by $a_{i}$. So, if $a_{k}$ is the agent to change state to Gather it must be before round $t^{\prime}+3n+1$. Thus $a_{k}$ must move into state Gather from state Forward at round $t^{\prime}+n+1$. This can only occur if $a_{k}$ finds a visited type message at $home$ of $a_{j}$ at round $t^{\prime}+n+1$. But this is not possible as $a_{j}$ have already erased all previous data on its $home$ at round $t^{\prime}$ while in state Initial, and there is no other agent which can visit $home$ of $a_{j}$ and alter the data between $t^{\prime}$ and $t^{\prime}+n+1$ rounds. Hence, at round $t^{\prime}+n+1$ when $a_{k}$ visits the $home$ of $a_{j}$ it does not find any visited type message. Thus $a_{i}$ must be the agent to change its state to Gather. Now to prove that $a_{i}$ stores the dir type message with direction component clockwise (which means $a_{j}$ was in state Forward while it was consumed), we have to show that $a_{i}$ changes its state to Gather from Backtrack. If possible let $a_{i}$ changes its state to Gather from state Forward. Then it must be at round $t^{\prime}+n+1$ when $a_{i}$ checks and finds a visited type message at the $home$ of $a_{k}$. Now at round $t^{\prime}$, $a_{k}$ which was in state Initial, cleared all previous data on its $home$. This implies the visited type message that $a_{i}$ finds at round $t^{\prime}+n+1$ must be written there after round $t^{\prime}$. But as only $a_{i}$ can be there after $t^{\prime}$ and before $t^{\prime}+n+1$, and since it can not alter any data on $home$ of $a_{k}$ before round $t^{\prime}+n+1$, it finds no visited type message at $home$ of $a_{k}$ at round $t^{\prime}+n+1$. Thus $a_{i}$ can not change its state to Gather from state Forward. Hence, it must change its state to Gather from state Backtrack at round $t^{\prime}+3n+1$ and according to the algorithm PerpExplore-Scat-WhitBrd, the dir type message that $a_{i}$ stores must have direction component clockwise.

### v2

@@ 91             Set $Move=0$
107                   Declare $Current-Node$ as the Byzantine black hole node and continue perpetual exploration avoiding this node.
@@
125             Declare the next node along the direction component of dir as the Byzantine black hole node and continue perpetual exploration avoiding this node.
@@
An agent can change its state to Gather2, only from the states Back-Wait or Gather. An agent changes its state to Gather2 from Gather, only when while moving along counter-clockwise direction finds another agent present, in which case it updates the dir message and writes the corresponding message of type dir and correspondingly changes its state to Gather2.
On the other hand, an agent can only change its state to Gather2 from Back-Wait, if it is waiting at its corresponding segments nearest clockwise $home$, and within which it encounters another agent at its current node along with it a message of type dir is written (in which the direction component of dir is along counter-clockwise). After storing the dir type message in its local memory and changing its state to Gather2 from Back-Wait, an agent immediately moves counter-clockwise in the same round. This generates another observation similar to the Observation 11.1 as follows.
{observation}
If two agents move to state Gather2, one moves from state Gather, and the other moves from state Back-Wait. Furthermore, they change to state Gather2 at the same node and leave the node together at the same round and stay together.
In state Gather2, the agent moves in a counter-clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID’)) type message, where ID’ matches with ID component of dir type message in its local memory. Note that our algorithm ensures that the current node is the nearest clockwise $home$ of $a_{1}$ if $a_{1}$ is the agent to be destroyed by the Byzantine black hole (refer Lemma 11.11). Further, as there are two agents at the current node, both in state Gather2 (refer Observation 11.1), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $left$, since the direction component of dir is counter-clockwise, and further changes its state to Cautious-Leader. On the other hand, the other alive agent at the current node, changes its state to Cautious-Follower, while initializing $Move=0$.
In state Cautious-Leader, an agent updates $Move=1$ and moves along the direction component of dir. After moving when it is alone on a node, it checks if the current node is marked with $Marking$ (i.e., either $left$ or $right$). If the node is marked, then it moves in the opposite direction (i.e., the opposite of direction component of dir message) to meet with the agent in state Cautious-Follower. Otherwise, if there is no $Marking$ at the current node then it identifies the current node as the Byzantine black hole node (only if it remains alive) and continues perpetual exploration avoiding this node. If an agent in the state Cautious-Leader with $Move=1$ meets with another agent (i.e., the agent in the state Cautious-Follower) it updates variable $Move$ to 0 and again moves according to the direction component.
In state Cautious-Follower, the agent after finding $Move=0$, initializes the variable $Wait=0$ and updates $Move=1$. If $Move=1$, then the agent checks whether $Wait<1$, if so then update $Wait=Wait+1$. When $Wait=1$, it checks if the current node has more than one agent (i.e., whether the agent in state Cautious-Leader has returned or not), if so then further update $Move=0$ and moves along the direction component of dir, along with the agent in state Cautious-Leader. Otherwise, if the agent in state Cautious-Leader does not return, that means the number of agents at the current node is 1 when $Wait=1$, this symbolizes that the agent in state Cautious-Leader has been either destroyed by the Byzantine black hole or the agent in state Cautious-Leader finds that its current node (i.e., the next node along the direction component of dir for the agent in state Cautious-Follower) is the Byzantine black hole and continues to perpetually move in the same direction, further avoiding this node. In any case, the agent executing Cautious-Leader has not returned implies the next node along the direction component of dir for the agent executing Cautious-Follower is the Byzantine black hole node. So, in this case, the agent declares the next node along the direction component of dir is the Byzantine black hole node and continues perpetual exploration avoiding that node.
11.2 Correctness and Complexity
In the next two lemmas (Lemma 11.1 and Lemma 11.3) we first ensure that when we say an agent encounters an anomaly, is actually an anomaly. And what an agents interpret from these anomalies are true. Next in Lemma 11.5 we ensure that the agent that changes to state Gather can always identify the state of the destroyed agent at the time it gets destroyed at the Byzantine black hole.
@@ 11.2 Correctness and Complexity
Let $a_{i}$ and $a_{j}$ be two agents exploring the segments $Seg(a_{i})$ and $Seg(a_{j})$, such that $Seg(a_{i})\cap Seg(a_{j})=v_{c}$, where $v_{c}$ is also the $home$ of $a_{j}$. If during the execution of PerpExplore-Scat-Whitbrd, there exists a round $t>0$, in which $a_{i}$ in state Forward finds a visited type message, then there exists a round $0<t^{\prime}<t$, in which $a_{j}$ has been destroyed by the Byzantine black hole while exploring the segment $Seg(a_{j})$ in state Backtrack.
Proof 11.2.
@@
Let us suppose, there does not exist any round $t^{\prime}<t$ in which $a_{j}$ has been destroyed by the Byzantine black hole in state Backtrack. This means that either $a_{j}$ is yet to be destroyed by the Byzantine black hole at round $t^{\prime}$, for all $t^{\prime}<t$ or $a_{j}$ has been destroyed by the Byzantine black hole at round $t^{\prime}$, while it is in state Forward.
@@
Case I: Let $a_{j}$ is yet to be destroyed by the Byzantine black hole at round $t^{\prime}$, for all $t^{\prime}<t$. Note that since $a_{i}$ is at $v_{c}$ and checks for the visited type message, this implies $T_{time}=n$. This means at $(t-n-1)$-th, round $a_{i}$ changed its state to Forward from Initial. This also implies $a_{j}$ was also in state Initial at $(t-n-1)-$th round at $v_{c}$. Hence, at this round, $a_{j}$ must have cleared any whiteboard data at $v_{c}$ (as $v_{c}$ is the $home$ of $a_{j}$) and also has moved clockwise after changing the state to Forward. Observe that, the node $v_{c}$ can only be explored by $a_{i}$ and $a_{j}$. So, these two agents have the possibility to write the visited type message at $v_{c}$ after round $t-n-1$ and before round $t$. Further, note that in the next $n+1$ rounds after $(t-n-1)-$th round (i.e., until round $t$), $a_{j}$ cannot return back to $v_{c}$. So, $a_{j}$ can not write anything on $v_{c}$ after round $t-n-1$ and before round $t$. Also, $a_{i}$ can only write the visited type message at $v_{c}$ at the $t-$th round (i.e., when $T_{time}=n$). But, in this case, $a_{i}$ at $t-$th round before writing a message, finds that there already exists a visited type message, and this a contradiction, to the fact that $a_{j}$ has not been destroyed by the Byzantine black hole at round $t^{\prime}$, for all $t^{\prime}<t$.
@@
Case II: Let $a_{j}$ has been destroyed by the Byzantine black hole at some round $t^{\prime}<t$, while in state Forward. This implies there exists a round $0<t^{\prime\prime}<t^{\prime}$ in which $a_{j}$ was in state Initial (this is the last time $a_{j}$ was in state Initial before it gets destroyed by the Byzantine black hole). Thus this means, that $t^{\prime}\leq t^{\prime\prime}+n-2$, as a segment can be of length at most $n-2$. Then at round $t^{\prime\prime}$, $a_{i}$ was also in state Initial at its own $home$. Now, note that after $t^{\prime\prime}$, $a_{i}$ can visit $v_{c}$ within round $t^{\prime\prime}+1$ and $t^{\prime\prime}+n+1$ while in state Forward. Note that from round $t^{\prime\prime}+1$ and before round $t^{\prime\prime}+n+1$ no agent can write a visited type message at $v_{c}$ (by a similar argument as in Case I). So during round $t^{\prime\prime}+n+1$, $a_{i}$ can not see any visited type message at $v_{c}$. So, $t>t^{\prime\prime}+n+1$. Also note that the next time after round $t^{\prime\prime}+n+1$, $a_{i}$ visits $v_{c}$ in state Forward, is earliest at round $t^{\prime\prime}+4n+1$. This implies $t\geq t^{\prime\prime}+4n+1$. Now, let us consider the agent $a_{k}$ where $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{k}$ and $Seg(a_{j})\cap Seg(a_{i})=home$ of $a_{i}$. Note that in round $t^{\prime\prime}+n+1$ both $a_{k}$ and $a_{i}$ move into state Back-Wait (as none of them sees any visited type message). Next after waiting there up to round $t^{\prime\prime}+2n$, both of them change their state to Backtrack at round $t^{\prime\prime}+2n+1$. Next, they reach their corresponding $home$ and check for a visited type message at round $t^{\prime\prime}+3n+1$. Note that $a_{i}$ finds the visited type message left by $a_{k}$ and changes to state Initial-Wait whereas, $a_{k}$ does not find any visited type message left by $a_{j}$ (as $a_{j}$ is destroyed at the Byzantine black hole at state Forward at round $t^{\prime}\leq t^{\prime\prime}+n-2<t^{\prime\prime}+3n+1$). So, $a_{k}$ changes to state Gather1 at round $t^{\prime\prime}+3n+1$ and moves clockwise with the message dir$=$ (Clockwise, NULL) until it finds $a_{i}$ at it’s $home$ in state Initial-Wait before round $t^{\prime\prime}+4n$ (as length of a segment can be of length at most $n-2$). Thus before $t^{\prime\prime}+4n+1$-th round $a_{i}$ changes its state to Gather1. Since no agent moves to state Forward from state Gather1, $a_{i}$ can not be at $v_{c}$ in state Forward during round $t$. So we again arrive at a contradiction. This implies $a_{j}$ can not be in state Forward while it is destroyed at the Byzantine black hole at round $t^{\prime}<t$.
@@
Let $a_{i}$ and $a_{j}$ be two agents such that $v_{c}=Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$. If there exists a round $t>0$ such that $a_{i}$ checks and finds no visited type message on $v_{c}$ while in state Backtrack, then there must be a round $t^{\prime}<t$ in which $a_{j}$ was destroyed by the Byzantine black hole while in state Forward.
Proof 11.4.
@@
Let us consider that there does not exist any round $t^{\prime}<t$ where $a_{j}$ was destroyed by the Byzantine black hole in the state Forward. This implies that either $a_{j}$ is not destroyed by the Byzantine black hole for all rounds $t^{\prime}<t$ or it is destroyed by the Byzantine black hole at round $t^{\prime}<t$ while it was in state Backtrack.
@@
Case-I: Let us consider that $a_{j}$ is yet to be destroyed by the Byzantine black hole at round $t^{\prime},\leavevmode\nobreak\ \forall t>t^{\prime}>0$. Note that, at round $t$, $a_{i}$ checks for a visited type message while in state Backtrack, and that means the current $T_{time}$ for any alive agent must be equal $3n$ (at $T_{time}=n$ every alive agent ends its state Forward, then until $T_{time}=2n$ every such agent is in state Back-Wait, and at $T_{time}=3n$ every alive agent checks for the message in state Backtrack). Note that, this means at time $t-3n-1$, both $a_{i}$ and $a_{j}$ were in state Initial at their respective $home$. Then at round $t-2n-1$, both of them must be in state Forward with $T_{time}=2n$, while $a_{j}$ is at $v_{c}$ and $a_{i}$ is at $home$ of $a_{k}$ ( $a_{k}$ is the third agent exploring $Seg(a_{k})$, such that $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$). So, at this point, both $a_{i}$ and $a_{j}$ must have written a visited type message at their respective current nodes (i.e., $v_{c}$ and $home$ of $a_{k}$). Further observe that, within $t-2n$ round to $t-n-1$ round $a_{j}$ remains at $v_{c}$ while in state Back-Wait, and according to our algorithm no agent in state Back-Wait can alter the already stored information at their current node, i.e., within these rounds, $a_{j}$ cannot erase the visited type message at $v_{c}$. Next, within $t-n$ to $t-1$ round, the only agent that can visit the node $v_{c}$ is $a_{i}$, while it is in state Backtrack. Note that, $a_{i}$ cannot alter any visited type message during state Backtrack. Thus, $a_{i}$ at round $t$ checks and finds a visited type message at $v_{c}$. This is a contradiction to the assumption that $a_{i}$ does not find any visited type message at round $t$. Thus, $a_{j}$ must have been destroyed by the Byzantine black hole at time $t^{\prime}$ for some $t^{\prime}<t$.
@@
Case-II: Suppose $a_{j}$ has been destroyed by the Byzantine black hole at some round $t^{\prime}<t$, while in state Backtrack. Let $t^{\prime\prime}$ be the round when $a_{j}$ was in state Initial for the last time where $0<t^{\prime\prime}<t^{\prime}$. Note that $t^{\prime}>t^{\prime\prime}+2n+1$ (because an agent changes its state to Backtrack at round $t^{\prime\prime}+2n+1$). So at the round, $t^{\prime\prime}+n+1$ it was alive and at $v_{c}$ in state Forward. During this round, it also writes a visited type message at $v_{c}$. Now as we discussed earlier in Case-I, from round $t^{\prime\prime}+n+1$ till round $t-1$ no agent can erase or alter the visited type message at $v_{c}$. Thus at round $t$, $a_{i}$ must find the visited type message at $v_{c}$ upon checking while in state Backtrack. This is a contradiction to the fact that at round $t$, $a_{i}$ checks and finds no visited type message at $v_{c}$ while it is in state Backtrack. Thus if $a_{j}$ is destroyed by the Byzantine black hole at round $t^{\prime}$, then it must be in state Forward.
@@
The agent which changes its state to Gather, correctly identifies the state in which an agent was in, just before it gets destroyed by the Byzantine black hole.
Proof 11.6.
@@
Let $a_{i},a_{j}$, and $a_{k}$ be three agents initially at three different nodes (i.e., three different $home$ for three agents) of $R$. Let $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$. Without loss of generality let $a_{j}$ be the agent that is destroyed by the Byzantine black hole at a round $t>0$. Now we have two cases according to the state of $a_{j}$ at the time of getting destroyed.
@@
Case-I: $a_{j}$ gets destroyed by the Byzantine black hole while in state Forward at round $t$. Let $t^{\prime}<t$ be the round when $a_{j}$ was in state Initial for the last time. Note that at round $t^{\prime}$, all agents are in state Initial. We claim that $a_{i}$ is the agent to change its state to Gather. If possible let $a_{k}$ change its state to Gather. Then first we show that it must change its state to Gather before round $t^{\prime}+3n+1$. Otherwise, since $a_{j}$, fails to reach $home$ of $a_{i}$ at round $t^{\prime}+n+1$, it does not write any visited type message there but $a_{i}$ reaches $home$ of $a_{k}$ at round $t^{\prime}+n+1$ and writes a visited type message at $home$ of $a_{k}$. Also, since, $a_{i}$ and $a_{j}$ are the only two agents which visit $home$ of $a_{i}$., this implies, no message can be written at $home$ of $a_{i}$ on and between rounds $t^{\prime}$ and $t^{\prime}+3n$. So, when $a_{i}$ returns $home$ and finds no visited type message round $t^{\prime}+3n+1$ while in state Backtrack, it changes its state to Gather but $a_{k}$ doesn’t do so as it sees the visited type message at its $home$, left by $a_{i}$. So, if $a_{k}$ is the agent to change state to Gather it must be before round $t^{\prime}+3n+1$. Thus $a_{k}$ must move into state Gather from state Forward at round $t^{\prime}+n+1$. This can only occur if $a_{k}$ finds a visited type message at $home$ of $a_{j}$ at round $t^{\prime}+n+1$. But this is not possible as $a_{j}$ have already erased all previous data on its $home$ at round $t^{\prime}$ while in state Initial, and there is no other agent which can visit $home$ of $a_{j}$ and alter the data between $t^{\prime}$ and $t^{\prime}+n+1$ rounds. Hence, at round $t^{\prime}+n+1$ when $a_{k}$ visits the $home$ of $a_{j}$ it does not find any visited type message. Thus $a_{i}$ must be the agent to change its state to Gather. Now to prove that $a_{i}$ stores the dir type message with direction component clockwise (which means $a_{j}$ was in state Forward while it was destroyed), we have to show that $a_{i}$ changes its state to Gather from Backtrack. If possible let $a_{i}$ change its state to Gather from state Forward. Then it must be at round $t^{\prime}+n+1$ when $a_{i}$ checks and finds a visited type message at the $home$ of $a_{k}$. Now at round $t^{\prime}$, $a_{k}$ which was in state Initial, cleared all previous data on its $home$. This implies the visited type message that $a_{i}$ finds at round $t^{\prime}+n+1$ must be written there, after round $t^{\prime}$. But as only $a_{i}$ can be there after $t^{\prime}$ and before $t^{\prime}+n+1$, and since it can not alter any data on $home$ of $a_{k}$ before round $t^{\prime}+n+1$, it finds no visited type message at $home$ of $a_{k}$ at round $t^{\prime}+n+1$. Thus $a_{i}$ can not change its state to Gather from state Forward. Hence, it must change its state to Gather from state Backtrack at round $t^{\prime}+3n+1$ and according to the algorithm PerpExplore-Scat-WhitBrd, the dir type message that $a_{i}$ stores must have direction component clockwise.